付録
7.3: Using the Central Limit Theorem7.3:中心極限定理の利用
88. What does the law of large numbers say about the relationship between the sample mean and the population mean?
88。標本平均と母平均の関係について、大数の法則は何を述べているか。
89. Applying the law of large numbers, which sample mean would expect to be closer to the population mean, a sample of size ten or a sample of size 100?
89。大数の法則によると、大きさ10と大きさ100の標本では、どちらの標本平均が母平均に近いと予想されるか。
Use this information for the next three questions. A manufacturer makes screws with a mean diameter of 0.15 cm (centimeters) and a range of 0.10 cm to 0.20 cm; within that range, the distribution is uniform.
英語のヒント
次の情報を用いて、続く3問に答えよ。あるメーカーは、平均直径0.15cm、直径の範囲0.10~0.20cmのねじを製造している。この範囲内では一様分布である。
90. If X = the diameter of one screw, what is the distribution of X?
90。X = ねじ1本の直径、とすると、X の分布は何か。
91. Suppose you repeatedly draw samples of size 100 and calculate their mean. Applying the central limit theorem, what is the distribution of these sample means?
91。大きさ100の標本を繰り返し抽出し、平均を計算するとする。中心極限定理を適用すると、標本平均の分布は何か。
92. Suppose you repeatedly draw samples of 60 and calculate their sum. Applying the central limit theorem, what is the distribution of these sample sums?
92。大きさ60の標本を繰り返し抽出し、和を計算するとする。中心極限定理を適用すると、標本の和の分布は何か。
Practice Test 2 Solutions模擬試験2の解答
Probability Distribution Function (PDF) for a Discrete Random Variable離散確率変数の確率分布関数(PDF)
1. The domain of X = {English, Mathematics,….], i.e., a list of all the majors offered at the university, plus “undeclared.”
1。X の取りうる値は {英語、数学、…}、すなわち、その大学にあるすべての専攻と「未定」である。
2. The domain of Y = {0, 1, 2, …}, i.e., the integers from 0 to the upper limit of classes allowed by the university.
2。Y の取りうる値は {0, 1, 2, …}、すなわち、0から大学が認める履修科目数の上限までの整数である。
3. The domain of Z = any amount of money from 0 upwards.
3。Z の取りうる値は、0以上の任意の金額である。
4. Because they can take any value within their domain, and their value for any particular case is not known until the survey is completed.
4。それぞれの範囲内でさまざまな値を取り、個々の回答での値は調査が終わるまでわからないからである。
5. No, because the domain of Z includes only positive numbers (you can’t spend a negative amount of money). Possibly the value –7 is a data entry error, or a special code to indicated that the student did not answer the question.
5。いいえ。Z の取りうる値には正の数だけが含まれるためである。負の金額を使うことはできない。−7 は入力誤りか、学生が回答しなかったことを示す特別なコードかもしれない。
6. The probabilities must sum to 1.0, and the probabilities of each event must be between 0 and 1, inclusive.
6。確率の合計は1.0でなければならず、各事象の確率は0以上1以下でなければならない。
7. Let X = the number of books checked out by a patron.
7。X = 利用者が借りる本の冊数、とする。
8. P(x > 2) = 0.10 + 0.05 = 0.15
8。P(x > 2) = 0.10 + 0.05 = 0.15
9. P(x ≥ 0) = 1 – 0.20 = 0.80
9。P(x ≥ 0) = 1 − 0.20 = 0.80
10. P(x ≤ 3) = 1 – 0.05 = 0.95
10。P(x ≤ 3) = 1 − 0.05 = 0.95
11. The probabilities would sum to 1.10, and the total probability in a distribution must always equal 1.0.
11。確率の合計が1.10になるが、分布の確率の合計は常に1.0でなければならない。
12. = 0(0.20) + 1(0.45) + 2(0.20) + 3(0.10) + 4(0.05) = 1.35
Mean or Expected Value and Standard Deviation平均(期待値)と標準偏差
13.
x |
P(x) |
xP(x) |
|---|---|---|
30 |
30/130 |
6.92 |
40 |
40/130 |
12.31 |
60 |
60/130 |
27.69 |
14. = 6.92 + 12.31 + 27.69 = 46.92
15. P(x = 30) = 0.23
P(x = 40) = 0.31
P(x = 60) = 0.46
P(x = 40) = 0.31
P(x = 60) = 0.46
15。P(x = 30) = 0.23、P(x = 40) = 0.31、P(x = 60) = 0.46。
16.
x |
P(x) |
xP(x) |
(x – μ)2P(x) |
|---|---|---|---|
30 |
0.23 |
6.92 |
(30 – 46.92)2(0.23) = 66.09 |
40 |
0.31 |
12.31 |
(40 – 46.92)2(0.31) = 14.75 |
60 |
0.46 |
46.92 |
(60 – 46.92)2(0.46) = 78.93 |
Binomial Distribution二項分布
18. q = 1 – 0.65 = 0.35
18。q = 1 − 0.65 = 0.35
19.
- There are a fixed number of trials.試行回数が固定されている。
- There are only two possible outcomes, and they add up to 1.結果は二つだけで、その確率の合計は1である。
- The trials are independent and conducted under identical conditions.試行は独立で、同じ条件のもとで行われる。
20. No, because there are not a fixed number of trials
20。いいえ。試行回数が固定されていないため。
21. X ~ B(100, 0.65)
21。X ~ B(100, 0.65)
22. μ = np = 100(0.65) = 65
22。μ = np = 100(0.65) = 65
23.
24. X = Joe gets a hit in one at-bat (in one occasion of his coming to bat)
24。X = ジョーが1打数(1回打席に立つ機会)で安打を打つこと。
25. X ~ B(20, 0.4)
25。X ~ B(20, 0.4)
26. μ = np = 20(0.4) = 8
26。μ = np = 20(0.4) = 8
27.
4.4: Geometric Distribution4.4:幾何分布
28.
- A series of Bernoulli trials are conducted until one is a success, and then the experiment stops.ベルヌーイ試行を成功が1回出るまで繰り返し、そこで実験を終える。
- At least one trial is conducted, but there is no upper limit to the number of trials.少なくとも1回は試行するが、試行回数に上限はない。
- The probability of success or failure is the same for each trial.成功と失敗の確率は、毎回同じである。
29. T T T T H
29。T T T T H
30. The domain of X = {1, 2, 3, 4, 5, ….n}. Because you are drawing with replacement, there is no upper bound to the number of draws that may be necessary.
30。X の取りうる値は {1, 2, 3, 4, 5, …, n}。復元抽出なので、必要となる抽出回数に上限はない。
31. The domain of X = {1, 2, 3, 4, 5, 6, 7, 8., 9, 10, 11, 12…27}. Because you are drawing without replacement, and 26 of the 52 cards are red, you have to draw a red card within the first 17 draws.
31。X の取りうる値は {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, …, 27}。非復元抽出で、52枚中26枚が赤いため、最初の17回以内に赤いカードを引くことになる。
32. X ~ G(0.24)
32。X ~ G(0.24)
33.
34.